# Deep Linear Hawkes Processes

### v1

@@ Abstract
Marked temporal point processes (MTPPs) are used to model sequences of different types of events with irregular arrival times, with broad applications ranging from healthcare and social networks to finance. We address shortcomings in existing point process models by drawing connections between modern deep state-space models (SSMs) and linear Hawkes processes (LHPs), culminating in an MTPP that we call the deep linear Hawkes process (DLHP). The DLHP modifies the linear differential equations in deep SSMs to be stochastic jump differential equations, akin to LHPs. After discretizing, the resulting recurrence can be implemented efficiently using a parallel scan. This brings parallelism and linear scaling to MTPP models. This contrasts with attention-based MTPPs, which scale quadratically, and RNN-based MTPPs, which do not parallelize across the sequence length. We show empirically that DLHPs match or outperform existing models across a broad range of metrics on eight real-world datasets. Our proposed DLHP model is the first instance of the unique architectural capabilities of SSMs being leveraged to construct a new class of MTPP models.
@@ 1 Introduction
Figure 1: Intensity estimates from trained models when conditioned on an empty sequence $\mathcal{H}_{t}=\emptyset$ for NHP (Mei & Eisner, 2017) and DLHP, our method. Shown in dotted lines are the ground truth, inhomogeneous Poisson process intensity. Our DLHP is able to accurately capture the background intensity. See Section 5.1 for more details.
@@
Marked temporal point processes (MTPPs) are used to model irregular sequences of events in continuous-time, where each event has an associated type, often referred to as a mark. MTPPs model the joint distribution of marked event sequences. They have been successfully applied to modeling purchasing patterns in e-commerce (Türkmen et al., 2019; Vassøy et al., 2019; Yang et al., 2018), patient-specific medical events (Hua et al., 2022), disease propagation (Gajardo & Müller, 2023), and many other domains (Williams et al., 2020; Sharma et al., 2018; Wang et al., 2024).
@@
An MTPP is fully characterized by a marked intensity process which specifies the expected instantaneous rate of occurrence of events of each mark conditioned on the event history. State-of-the-art methods use neural networks to compute hidden states that summarize the event history, which are then used to compute marked intensities across future values of time. However, many models are limited by inexpressive temporal dynamics, lack of support for long-range dependencies, and serial computation (Du et al., 2016; Mei & Eisner, 2017). Recent advances in transformer-based MTPPs have improved performance and gained parallelism, but scale quadratically with sequence lengths (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022).
@@
Recently, deep state-space models (often abbreviated as SSMs) have emerged as a challenger to transformer-based models for discrete sequence modeling (Gu et al., 2022b; Smith et al., 2022; Gu & Dao, 2023). SSMs interleave a stack of linear state-space recurrences with position-wise non-linearities (Gu et al., 2021). This architecture has been found to be not only highly performant on a wide range of tasks (Goel et al., 2022; Deng et al., 2024), but retains linear scaling, can be parallelized across the length of a sequence, and can gracefully handle irregularly-spaced observations.
Inspired by this, we revisit a foundational point process model, the linear Hawkes process (LHP Hawkes, 1971), and draw connections between LHPs and deep SSMs. We combine the parameterization and parallelization strategy of SSMs with the functional form of LHPs to create what we call the deep linear Hawkes process (DLHP). More formally, the DLHP is a fully-recurrent neural MTPP parameterized by a stack of stochastic jump differential equations on the complex plane (serving as the recurrence) interleaved with position-wise non-linearities (to improve expressivity). This design yields an MTPP with two main advantages over existing neural MTPPs: (i) parallelism across the length of the sequence through the use of parallel scans, and (ii) highly flexible intensity functions. This is achieved not only through the expressivity of the SSM-style architecture, but also by tying the output intensity at time $t$ to the model’s continuously-evolved hidden state $\mathbf{x}_{t}$ (extending ideas from Mei & Eisner (2017) and Yang et al. (2022), see Figs. 1 and 2), and by going beyond the classical LHP form with input-dependent recurrent dynamics (akin to Mamba (Gu & Dao, 2023)).
The contributions of this paper are as follows: We introduce a new family of marked point process models, deep linear Hawkes processes—the first MTPP model that fully leverages the architectural features of deep SSMs. We demonstrate that DLHPs match or exceed the performance of existing models across eight real-world datasets, with an average per-event likelihood improvement of 38% across datasets, over the individually best-performing existing method per dataset. We also verify that DLHP scales more effectively to longer sequences, a crucial capability for a wide range of applications. We will release our models, datasets, and pipelines as part of the existing EasyTPP library (Xue et al., 2023). We conclude by discussing the relative advantages and disadvantages of the DLHP over existing methods and opportunities for extending this work.
Figure 2: Three different schematics of the deep linear Hawkes process (DLHP) and latent linear Hawkes (LLH) layer we propose. With increasing granularity: Left (a): On a high level, the DLHP can simply be viewed as a deep stack of neural network layers that transform an event sequence into an intensity function. Middle (b): On a more granular level, individual LLH layers can be viewed as discrete-time recurrences (see Eq. 16), directly defining an intensity evaluated at select times: $t^{*}$ using $\mathbf{x}_{t^{*}}$, right limits $t_{i}+$ using $\mathbf{x}_{t_{i}}$, and left values $t_{i}$ using $\mathbf{x}_{t_{i}-}$. Right (c): The same recurrences can be viewed as a set of non-linearly coupled stochastic jump differential equations in continuous-time. Events are embedded and impart impulses to the differential equation. Impulses for layer $L$ are omitted for visual clarity. Intermediate intensities are evaluated using $\vec{0}$ for $\mathbf{u}^{(1)}$ and no impulse.
@@ 2.1 Marked Temporal Point Processes
Let $t_{1},t_{2},\dots\in\mathbb{R}_{\geq 0}$ be a strictly increasing sequence of positive random variables, each representing the time of occurrence for an event of interest. [^1] Please refer to Tables 2 and 3 in Appendix A for a list of notation and acronym definitions, respectively. For each $t_{i}$, let $k_{i}\in\mathcal{M}$ be a random variable representing accompanying side-information, commonly referred to as an event’s mark, with $\mathcal{M}$ being the mark-space. In this paper, we focus on discrete and finite mark spaces, i.e., :=$\mathcal{M}:=\{1,\dots,K\}$; however, in general $\mathcal{M}$ can be continuous or even a mixture of continuous and discrete. Together $t_{i}$ and $k_{i}$ fully define a given event. The joint distribution over a sequence of continuous event times and mark types is described as a marked temporal point process. We use $\mathcal{H}_{t}$ to represent the sequence, or history, of events up to some time $t$: $\mathcal{H}_{t}:=\{(t_{i},k_{i})\;|\;t_{i}\leq t\text{ for }i\in\mathbb{N}\}$, with $\mathcal{H}_{t-}$ defined similarly except that it does not include events that occur at time $t$.
@@
One way of characterizing an MTPP is through a marked intensity process, which describes the instantaneous expected rate of occurrence for events of specific marks. Let $\mathbf{N}_{t}:=[N_{t}^{1},\dots,N_{t}^{K}]^{\top}\in\mathbb{Z}_{\geq 0}^{K}$ be the marked counting process which represents the number of occurrences of events of each type of mark in the time span $[0,t]$. The marked intensity process $\bm{\lambda}_{t}:=[\lambda_{t}^{1},\dots,\lambda_{t}^{K}]^{\top}\in\mathbb{R}_{\geq 0}^{K}$ characterizes an MTPP by describing how the counting process changes via:
@@
$\lambda_{t}^{k}\mathrm{d}t:=$ $\;\mathbb{E}\left[\text{event of type }k\text{ occurs in }[t,t+\mathrm{d}t]\;|\;\mathcal{H}_{t-}\right]=\;\mathbb{E}\left[N^{k}_{t+\mathrm{d}t}-N^{k}_{t}\;|\;\mathcal{H}_{t-}\right],$ (1)
with the total intensity $\lambda_{t}:=\sum_{k=1}^{K}\lambda^{k}_{t}$ being the rate that any event occurs. Note that the intensity conditions on the left limit of the history $\mathcal{H}_{t-}$ to ensure that the intensity is modeling future events.
@@
Parameterized forms of $\bm{\lambda}$ are often trained by optimizing the log-likelihood over observed data. The log-likelihood for a single sequence $\mathcal{H}_{T}$ is defined as (Daley & Vere-Jones, 2003, ch. 7.3):
$\mathcal{L}(\mathcal{H}_{T}):=\sum\nolimits_{i=1}^{|\mathcal{H}_{T}|}\log\lambda_{t_{i}}^{k_{i}}-\int_{0}^{T}\lambda_{s}\mathrm{d}s.$ (2)
Linear Hawkes Processes  An (unmarked) Hawkes process (Hawkes, 1971), or more generally a self-exciting process, is a temporal point process where event occurrences increase the rate at which subsequent events occur soon thereafter. Of particular interest to us are linear Hawkes processes (LHPs), which are characterized by the following intensity process:
$\lambda_{t}:=\nu+\int_{s=0}^{t-}h(t-s)\mathrm{d}N_{s}:=\nu+\sum\nolimits_{i=1}^{N_{t-}}h(t-t_{i}),$ (3)
where $\nu>0$ is the background intensity, $h:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is the excitation function (or kernel), and $N_{t}$ is the associated counting process characterized by intensity $\lambda_{t}$. $N_{t-}$ is used as the upper limit in Eq. 3 to ensure the intensity at time $t$ does not take into account an event that occurs at time $t$.
Should $h$ correspond to the exponential decay kernel, $h(z)=\alpha\exp\left(-\beta z\right)$, then the LHP intensity process is Markov (Law & Viens, 2016) and admits the following stochastic differential form:
@@
$\mathrm{d}\lambda_{t}=\beta(\nu-\lambda_{t-})\mathrm{d}t+\alpha\mathrm{d}N_{t}\iff\lambda_{t}$ $=\nu+\int_{0}^{t-}\alpha\exp\left(-\beta(t-s)\right)\mathrm{d}N_{s}$ (4)
$=\nu+\sum\nolimits_{i=1}^{N_{t-}}\alpha\exp\left(-\beta(t-t_{i})\right).$ (5)
LHPs can be extended to the marked setting, with $K$ possible discrete marks, by replacing $\nu$ with a vector of $K$ background rates $\bm{\nu}:=[\nu_{1},\dots,\nu_{K}]^{\top}$, and the excitation effect $h(t-s)\mathrm{d}N_{s}$ with $\mathbf{h}(t-s)\mathrm{d}\mathbf{N}_{s}$. Here, $\mathbf{h}_{ij}$ of $\mathbf{h}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{K\times K}$ describes the excitation that events of type $i$ exerts on future events of type $j$. The counting process, $\mathrm{d}\mathbf{N}_{t}$, is then either a $K$-dimensional zero-vector if no event occurs at time $t$, or a one-hot vector indicating which mark is associated with the occurring event. Generalizing the exponential kernel to handle marks results in the following differential form:
$\mathrm{d}\bm{\lambda}_{t}=-\bm{\beta}(\bm{\lambda}_{t-}-\bm{\nu})\mathrm{d}t+\bm{\alpha}\mathrm{d}\mathbf{N}_{t},$ (6)
where $\bm{\beta,\alpha}\in\mathbb{R}^{K\times K}_{\geq 0}$ are restricted to be non-negative to ensure non-negative marked intensities.
2.2 State-Space Models
@@
Deep state-space models (SSMs) are a recent innovation in recurrent models that have found success in long-range sequence modeling tasks (Gu et al., 2022b) and language modeling tasks (Gu & Dao, 2023), while also having favorable computational properties. The backbone of deep SSMs is the linear state-space equations, which define a continuous-time dynamical system with input and output signals $\textbf{u}(t),\mathbf{y}(t)\in\mathbb{R}^{H}$, respectively, through linear differential equations:
$\frac{\mathrm{d}}{\mathrm{d}t}\textbf{x}(t)$ $=\mathbf{A}\textbf{x}(t)+\mathbf{B}\mathbf{u}(t)$ (7)
$\mathbf{y}(t)$ $=\mathbf{C}\mathbf{x}(t)+\mathbf{D}\mathbf{u}(t),$ (8)
where $\textbf{x}(t)\in\mathbb{R}^{P}$ is the (hidden) state of the system, and $\mathbf{A}\in\mathbb{R}^{P\times P},\mathbf{B}\in\mathbb{R}^{P\times H},\mathbf{C}\in\mathbb{R}^{H\times P},$ and $\mathbf{D}\in\mathbb{R}^{H\times H}$ are the parameters that control the system’s dynamics.
@@
Deep SSMs then stack these recurrences interleaved with non-linear position-wise functions, $\sigma$. The function $\sigma$ can contain activation functions, residual connections and normalization layers, and transforms the output $\mathbf{y}$ of the previous recurrence into the input $\mathbf{u}$ of the next, i.e., $\mathbf{u}^{(l)}(t):=\sigma(\mathbf{y}^{(l-1)}(t))$ for layer $l$. This combination yields a sequence model where each recurrence is conditionally linear in time given the input, but is ultimately non-linear in depth due to the function $\sigma$.
To evaluate the SSM, we first discretize the continuous-time system at the desired times, and then evaluate as though it were a conventional discrete-time RNN architecture. Crucially, the linearity of the resulting discrete-time recurrence allows it to be evaluated using parallel scans (Blelloch, 1990; Smith et al., 2022; Gu & Dao, 2023), leading to linear work scaling (i.e., number of operations), and, importantly, sublinear scaling of the computation time with respect to sequence length given sufficient parallel compute. Note this contrasts with conventional sequential RNNs (e.g., LSTMs), which process sequences serially; and attention-based methods, which can be parallelized over a sequence, but have quadratic work scaling with respect to sequence length. This allows SSMs to fully and efficiently leverage modern massively parallel hardware while also a being highly expressive and performant model class. Importantly for our purposes, evaluating a linear recurrence with a parallel scan natively admits evaluations with varying observation intervals. We will leverage this to parsimoniously handle the variable inter-event times observed in MTPP settings.
3 Deep Linear Hawkes Processes
In this section, we introduce our deep linear Hawkes process (DLHP), a neural MTPP that draws a novel connection between LHPs and deep SSMs. Stochastic jump differential equations, akin to the LHP intensity, form the basis of the conditionally linear recurrent layer, which we refer to as a latent linear Hawkes (LLH) layer. The LLH layer can be viewed as a modified SSM recurrence, while still admitting parallel computation. Taking further inspiration from deep SSMs, the DLHP is then made up of a stack of LLH layers, each interleaved with non-linear, position-wise transformations to increase the overall expressivity of the model (see Fig. 2). In this section, we formalize this approach and outline implementation details.

### v2

# Deep Continuous-Time State-Space Models for Marked Event Sequences

@@ Abstract
Marked temporal point processes (MTPPs) model sequences of events occurring at irregular time intervals, with wide-ranging applications in fields such as healthcare, finance and social networks. We propose the state-space point process (S2P2) model, a novel and performant model that leverages techniques derived for modern deep state-space models (SSMs) to overcome limitations of existing MTPP models, while simultaneously imbuing strong inductive biases for continuous-time event sequences that other discrete sequence models (i.e., RNNs, transformers) do not capture. Inspired by the classical linear Hawkes processes, we propose an architecture that interleaves stochastic jump differential equations with nonlinearities to create a highly expressive intensity-based MTPP model, without the need for restrictive parametric assumptions for the intensity. Our approach enables efficient training and inference with a parallel scan, bringing linear complexity and sublinear scaling while retaining expressivity to MTPPs. Empirically, S2P2 achieves state-of-the-art predictive likelihoods across eight real-world datasets, delivering an average improvement of 33% over the best existing approaches.
@@ 1 Introduction
Marked temporal point processes (MTPPs) are used to model irregular sequences of events in continuous time, where each event has an associated type, often called a mark. MTPPs model the joint distribution of these sequences of event times and marks. They have been successfully applied to modeling purchasing patterns in e-commerce (Türkmen et al., 2019; Vassøy et al., 2019; Yang et al., 2018), patient-specific medical events (Hua et al., 2022), disease propagation (Gajardo and Müller, 2023), and event modeling and prediction across multiple other domains (Williams et al., 2020; Sharma et al., 2018; Wang et al., 2024). An MTPP can be fully characterized by a marked intensity process, specifying the instantaneous rate of occurrence of each mark conditioned on history.
@@
State-of-the-art neural methods compute hidden states to summarize the event history, which are then used to compute marked intensities at any point in time. However, many models are limited by inexpressive temporal dynamics, lack of support for long-range dependencies, and serial computation (Du et al., 2016; Mei and Eisner, 2017). Recent advances in transformer-based MTPPs improved performance and gained parallelism, but scale quadratically in sequence length (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022), preventing them from being used in practice for long sequences, such as modeling clinical events over a patient’s entire medical history spanning many years.
Rank (Lower ↓ is better, x = best, x = second-best)
Likelihood Next Event Prediction Calibration Composite
Model Mark Time Mark Acc. Time RMSE Mark Time Rank
RMTPP (Du et al., 2016) 6.8 6.1 6.9 5.0 5.8 6.1 6.1
SAHP (Zhang et al., 2020) 6.9 4.9 7.1 4.3 6.8 5.4 5.5
THP (Zuo et al., 2020) 5.5 7.0 5.6 4.3 5.0 5.4 5.5
IFTPP (Shchur et al., 2020a) 4.1 3.1 5.0 4.0 1.8 2.6 3.6
MHP (Gao et al., 2024) 5.4 6.4 4.3 5.5 4.8 6.0 5.3
NHP (Mei and Eisner, 2017) 2.4 3.3 1.8 2.4 4.9 3.6 2.9
AttNHP Yang et al. (2022) 2.4 3.3 2.9 6.6 3.4 3.7 3.7
S2P2 (Ours) 1.9 1.4 1.9 2.3 3.0 2.8 2.1
Figure 1: (a) A schematic of our proposed state-space point process (S2P2), a deep stack of novel latent linear Hawkes (LLH) layers interleaved with nonlinear and normalization layers creating an expressive MTPP architecture. (b) Summary table of results we present in Section˜6. We summarize ranks for six key metrics, ranking the average held-out test set performance across eight real-world datasets for five randomly seeded models; as well as a holistic composite rank, defined as the average of the ranks. Our S2P2 model outperforms all baselines by almost an entire rank, strongly indicating state-of-the-art performance and robustness across metrics and datasets.
@@
Recently, deep state-space models (often abbreviated as SSMs) have emerged as a challenger to transformer-based models for discrete sequence modeling (Gu et al., 2022a; Smith et al., 2022; Gu and Dao, 2023). SSMs interleave a stack of linear state-space recurrences with position-wise nonlinearities (Gu et al., 2021). This architecture not only achieves superior performance on a wide range of tasks (Goel et al., 2022; Deng et al., 2024), but retains linear scaling, can be parallelized across the length of a sequence, and can gracefully handle irregularly spaced observations.
Despite being defined in continuous-time, applying SSMs to MTPPs is not straightforward. Tying the conditional intensity to the model’s state is the obvious choice, but SSMs expect a continuous-valued signal to integrate as input, whereas event sequences are discontinuous by nature. Taking inspiration from classical parametric linear Hawkes processes (LHPs) (Hawkes, 1971), we introduce a stochastic jump differential equation on the complex plane to serve as the SSM recurrence. We refer to this as a latent linear Hawkes (LLH) layer. We derive closed-form update rules for any time that allow the diagonalized continuous-time SSM system to use efficient parallel scans. We then introduce our state-space point process (S2P2) model, a stack of LLH layers interleaved with position-wise nonlinear functions (Fig.˜1(a)), inspired by deep SSMs. This design yields an MTPP that has both a highly flexible conditional intensity function and access to efficient, parallelizable computation.
This paper is organized as follows: In Section˜2 we present the necessary preliminaries on MTPPs, LHPs, and deep SSMs. In Section˜3 we introduce our state-space point process architecture. We first build the connection between LHPs and deep SSMs, and then derive how to use this connection to make a highly expressive, parsimonious and parallelizable MTPP. In Section˜5 we then verify this expressivity on small-scale, targeted synthetic explorations comparing to key baselines. In Section˜6 we empirically evaluate our model at scale on a range of metrics across eight real-world datasets, finding that S2P2 matches or exceeds the average predictive performance of baselines, achieving either best- or second-best average performance on all six metrics. These results are summarized in Table 1(b). In Sections˜7 and 8 we conclude by discussing the relative advantages and limitations of S2P2 and promising future opportunities. Due to its robustness, performance, and efficiency, our S2P2 model is a powerful model out-of-the-box for a wide range of MTPP applications.
@@ 2.1 Marked Temporal Point Processes
We define an event sequence, or history, as a sequence of time-mark pairs, $\mathcal{H}_{t}:=\{(t_{i},k_{i})\;|\;t_{i}\leq t\text{ for }i\in\mathbb{Z}^{+}\}$, where $t_{i}\in\mathbb{R}_{\geq 0}$, $\forall i:t_{i-1}<t_{i}$, and $k_{i}\in\mathcal{M}$. [^1] Please refer to Tables 3 and 4 in Appendix A for a list of notation and acronym definitions, respectively. In this paper, we focus on discrete and finite mark spaces, i.e., $\mathcal{M}:=\{1,\dots,K\}$; however, $\mathcal{M}$ can be more general, such as countable or continuous. We also define $\mathcal{H}_{t-}$ similarly to $\mathcal{H}_{t}$, except that it excludes events at exactly time $t$.
@@
One way of characterizing an MTPP is through a marked intensity process. The intensity $\bm{\lambda}_{t}:=[\lambda_{t}^{1},\dots,\lambda_{t}^{K}]^{\top}\in\mathbb{R}_{\geq 0}^{K}$ characterizes an MTPP by describing the rate of occurance of events:
@@
$\lambda_{t}^{k}\mathrm{d}t:=\;\mathbb{E}\left[\text{event of type }k\text{ occurs in }[t,t+\mathrm{d}t)\;|\;\mathcal{H}_{t-}\right]$ (1)
with the total intensity $\lambda_{t}:=\sum_{k=1}^{K}\lambda^{k}_{t}$ being the rate that any event occurs. This intensity also defines a marked counting process $\mathbf{N}_{t}:=[N_{t}^{1},\dots,N_{t}^{K}]^{\top}\in\mathbb{Z}_{\geq 0}^{K}$, representing the number of occurrences of events of each type of mark in the time span $[0,t]$.
@@
Parameterized forms of $\bm{\lambda}$ are often trained by optimizing the log-likelihood over observed data (e.g., (Mei and Eisner, 2017; Zuo et al., 2020)). The log-likelihood for a single sequence $\mathcal{H}_{T}$ is defined as (Daley and Vere-Jones, 2003, ch. 7.3):
$\mathcal{L}(\mathcal{H}_{T}):=\sum\nolimits_{i=1}^{N_{T}}\log\lambda_{t_{i}}^{k_{i}}-\int_{0}^{T}\lambda_{s}\mathrm{d}s.$ (2)
Linear Hawkes Processes  A foundational MTPP is the linear Hawkes process (LHP). The LHP is a self-exciting process, where event occurrence increases the rate of occurrence of other events, with the influence decaying according to a kernel. The intensity function is the summation of influences, and if the kernel is the exponential function, it has the following integral and differential forms:
@@
$\bm{\lambda}_{t}=\bm{\nu}+\sum\nolimits_{i=1}^{N_{t-}}\exp\left(-\bm{\beta}(t-t_{i})\right)\bm{\alpha}\iff\mathrm{d}\bm{\lambda}_{t}=-\bm{\beta}(\bm{\lambda}_{t-}-\bm{\nu})\mathrm{d}t+\bm{\alpha}\mathrm{d}\mathbf{N}_{t},$ (3)
where, to ensure non-negative marked intensities, $\bm{\nu}\in\mathbb{R}^{K}_{\geq 0}$ and $\bm{\alpha},\exp{(-\bm{\beta})}\in\mathbb{R}^{K\times K}_{\geq 0}$ with $\exp$ being the matrix exponential. This form (which is the most common) of the LHP is incredibly limited, and hence is used primarily for its interpretability, as opposed to outright predictive performance.
2.2 Deep State-Space Models
@@
Deep state-space models (SSMs) are a class of recurrent models that have excelled in long-range sequence and language modeling tasks, all while having favorable computational properties (Gu et al., 2022a; Smith et al., 2022; Gu and Dao, 2023). The backbone of deep SSMs are linear state-space models, which define a continuous-time dynamical system with inputs and outputs $\textbf{u}(t),\mathbf{y}(t)\in\mathbb{R}^{H}$ through linear differential equations:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{x}(t)$ $=\mathbf{A}\mathbf{x}(t)+\mathbf{B}\mathbf{u}(t),\quad\mathbf{A}\in\mathbb{R}^{P\times P},\mathbf{B}\in\mathbb{R}^{P\times H}$ (4)
$\mathbf{y}(t)$ $=\mathbf{C}\mathbf{x}(t)+\mathbf{D}\mathbf{u}(t),\quad\mathbf{C}\in\mathbb{R}^{H\times P},\mathbf{D}\in\mathbb{R}^{H\times H}$ (5)
@@
where $\mathbf{x}(t)\in\mathbb{R}^{P}$ is the (hidden) state of the system, and $\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D}$ define the system’s dynamics. Deep SSMs interleave these recurrences with nonlinear position-wise functions $\sigma$ as $\mathbf{u}^{(l)}(t):=\sigma(\mathbf{y}^{(l-1)}(t))$ (for layer $l$). This yields a sequence model where each recurrence is conditionally linear in time but is nonlinear overall due to $\sigma$.
To evaluate an SSM, we first discretize the continuous-time system at appropriate times to yield a discrete sequence of closed-form state updates (Smith et al., 2022). The resulting discrete-time recurrence can be evaluated using parallel scans (Blelloch, 1990), with linear work scaling and, importantly, sublinear (theoretically logarithmic) scaling of the computation time with respect to sequence length given sufficient parallel compute. This parallel evaluation natively allows varying observation intervals or latent dynamics.
3 State-Space Point Process
We seek to define an MTPP model that is (a) highly expressive and (b) can access efficient and parallelizable compute methods. In this section, we formally introduce our state-space point process (S2P2) model, as outlined below: Section˜3.1 extends and generalizes the continuous-time form of the LHP, creating a layer we refer to as the latent linear Hawkes (LLH) layer. This generalization decouples the choice of the width of the layer from the mark space, allowing us to make arbitrarily wide layers, and is the basis for the novel connection between the LHP and deep SSMs. Section˜3.2 discusses how we make each LLH efficient and tractable to compute with parallel scans, and Section˜3.3 extends the model with time-varying dynamics to be more expressive. We conclude in Section˜3.4 by describing S2P2, a highly expressive MTPP model built from the composition of multiple LLH layers and nonlinearities that can be computed with a novel parallel inference scheme.